OpenStudy (anonymous):

What is the gravitational attraction between you and the Earth if you are at 7.5 x 10^6 m from the center of the Earth and your mass is 90 kg?

OpenStudy (anonymous):

@Data_LG2

OpenStudy (perl):

lets use newtons gravitational attraction formula $$ \Large F = \frac{GmM}{r^2}$$

OpenStudy (perl):

what is earths mass, and what is G ?

OpenStudy (anonymous):

whoops i forgot to add this but the mass of the earis is 5.98x10^24 kg

OpenStudy (perl):

ok thanks, and is G given, the universal gravity constant?

OpenStudy (anonymous):

nope but i feel like it would be 9.8 m/s

OpenStudy (perl):

nope, thats the acceleration of gravity near the surface of the earth

OpenStudy (perl):

we can look up 'universal gravitational constant'

OpenStudy (anonymous):

6.67x10^-11 m^3

OpenStudy (perl):

http://prntscr.com/76emgj

OpenStudy (anonymous):

whats next?

OpenStudy (perl):

$$ \Large F = \frac{GmM}{r^2} \\\Large F = \frac{6.67384 \times 10^{-11} \cdot 90 }{(7.5 \times 10^6)^2} $$

OpenStudy (anonymous):

then do i just simplify

OpenStudy (perl):

yes, i would use a calculator at this point

OpenStudy (anonymous):

okay! thanks so much for the help!!